OpenStudy (anonymous):

Evaluate the limit: lim h -> 0 (((5+h)^2 + 1)/(5+h) - 26/5 ) / h

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} (((5+h)^{2} + 1 / (5+h) ) - 26/5) / h\]

hero (hero):

Simplify, as much as possible, then substitute h = 0 into the expression.

hero (hero):

You want to simplify until h can cancel in the denominator.

hero (hero):

Once you have done that, then you can substitute h = 0 and find the limit

hero (hero):

I wasn't trying to butt in merstj. You can continue posting your solution. Just giving general steps

OpenStudy (anonymous):

no limit,

OpenStudy (anonymous):

i know... I'm having trouble simplifying...

OpenStudy (mertsj):

\[\frac{25+10h+h^2+\frac{1}{5+h}-\frac{26}{5}}{h}\]

OpenStudy (mertsj):

The limit does not exist because as you approach from the left the limit is negative infinity and as you approach from the right the limit is positive infinity